In a chess board one wheat grain is kept in the first box, two wheat grains are kept in the sec3nd uox, simiirt lr *h"rt gr;ii'i, .oded to the number of wheat grains in the pr.uiou, uol.'rh"n f,,rd il;iar number of wheat grains kept in aff the boxes irii..i.ss board

Grade:10

1 Answers

Sridharan N
7 years ago
This forms a geometric series with a=1, r=2, n=64(no.of squares),So total no. Of grains= 1(2^64-1)/2-1=2^64-1Best regards!